OpenStudy (lina17):

4(2x-1)-3=9(x-4) help?

OpenStudy (anonymous):

4(2x-1)-3 = 9(x-4) 8x - 4 - 3 = 9x - 36 8x - 7 = 9x - 36 8x - 7 - 9x + 36 = 0 x + 29 = 0 x = -29

OpenStudy (priyar):

@akemitan we should help the person arrive at the answer and not just give them the answer..coz that way they can do any problem of the same sort..

OpenStudy (anonymous):

@priyar im new here, will try to help next time

OpenStudy (priyar):

its ok...no problem..